OpenStudy (anonymous):

Multiply (6 – 7i)(3 – 6i)

OpenStudy (anonymous):

FOIL them out!

OpenStudy (anonymous):

@Chlorophyll He speaks the truth!

OpenStudy (anonymous):

ok, 18 + 36i - 21i + 42i

OpenStudy (anonymous):

now what?

OpenStudy (anonymous):

-24-57i

OpenStudy (anonymous):

Simplify them, will you!

OpenStudy (anonymous):

You foiled wrong!!

OpenStudy (anonymous):

-7i *– 6i= -7*-6*i*i= -42 The nagative cancels out and you got the 42 right but if you mutltiply two i you get -1

OpenStudy (anonymous):

18 - 42 i^2 - 36i - 21 i

OpenStudy (anonymous):

Can you simplify them now?

OpenStudy (anonymous):

Keep in mind i^2 = -1 :)

OpenStudy (anonymous):

-42i^2= 42

OpenStudy (anonymous):

No @Chlorophyll 18 - 42 i^2 - 36i - 21 i thahts wrong It should be 18 + 42 i^2 - 36i - 21 i

OpenStudy (anonymous):

I mean - 24 - 57i

OpenStudy (anonymous):

If you simplify them, that the end result -24 - 57i :)

OpenStudy (anonymous):

@romeo, thank for reminding, my mind just calculate i^2 too early :/